OpenStudy (anonymous):

PT=4/5TQ, PQ=18 Find PT. Segment Addition Postulate PT+TQ=PQ

OpenStudy (paxpolaris):

\[\large PT+TQ=PQ\]\[ \large \implies PT+TQ=18\] and since, \[\large PT=\frac 45 \times TQ\]\[\large \implies TQ=\underline{\frac 54 \times PT}\] replace this for TQ in the previous equation and solve for PT 

OpenStudy (paxpolaris):

\[\Large \implies PT + \frac 54 \cdot PT =18\]

OpenStudy (anonymous):

Why is it 5 over 4?

OpenStudy (paxpolaris):

if PT = 4/5 TQ , then TQ = 5/4 PT

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So multiply 18 times 5/4?

OpenStudy (paxpolaris):

first add PT + 5/4 PT

OpenStudy (anonymous):

Can you do the equation step by step to the answer because im confused

OpenStudy (paxpolaris):

\[PT +\frac 54PT \\= PT \left( 1+\frac54 \right)\]

OpenStudy (paxpolaris):

\[=PT \times \left( \frac94 \right)\]

OpenStudy (anonymous):

k got that whats next?

OpenStudy (paxpolaris):

multiply both sides by 4/9... \[PT \times \frac 94 \times \frac 49=18 \times \frac 49\] \[\implies PT \times \frac {\cancel9}4 \times \frac 4{\cancel9}=18 \times \frac 49\]\[\implies PT \times \frac {\cancel9}{\cancel4} \times \frac {\cancel4}{\cancel9}=18 \times \frac 49\] \[\Large \implies PT =18 \times \frac 49\]

OpenStudy (paxpolaris):

\[\Large \implies PT =2\cancel{18} \times \frac 4{\cancel9}=\color{green}8\]

OpenStudy (anonymous):

Thanks I understand so much more clearly now.